# Gaussian consensus processes and their Lyapunov exponents

### v1

@@ 2.2.2 The maximal Lyapunov exponent of $P(t)=M(t)\dots M(1)$ for $t\geq 0$
Following Newman [20], note that for our random matrix $M$, for any orthogonal matrix $Q\in O(N-1)$, the law of $Q^{-1}MQ=Q^{T}MQ$ coincides with the law of $M$. (One just has to check that the entries of $Q^{T}MQ$ are pairwise uncorrelated Gaussians with the correct means and variances, a simple calculation.) It follows that the invariant measure $\pi$ is also invariant with respect to the group $O(N-1)$. There is a unique such measure on $P\left(\mathbb{R}^{N-1}\right)$. It also follows that the distribution of $\frac{\|Mv\|}{\|v\|}$ does not depend on $v$, so
@@
where $\psi(x)=\frac{\Gamma^{\prime}(x)}{\Gamma(x)}$ is the digamma function. This follows from the statement that each non-central chi-squared random variable is a Poisson-weighted mixture of standard chi-squared random variables. Precisely, if $J\sim\mathrm{Poi}(\kappa/2)$ and, conditional on $J=j$, $Y$ is a $\chi^{2}$ r.v. with $\nu+2j$ degrees of freedom, then $Y\sim\chi^{2}_{\nu}(\kappa)$.
@@
$\begin{split}\lambda_{1}&=\frac{1}{2}\left[\log\frac{2\rho}{N}+\exp\left(-\frac{N\beta^{2}}{2\rho}\right)\,\sum_{j=0}^{\infty}\frac{\left(\frac{N\beta^{2}}{2\rho}\right)^{j}}{j!}\psi\left(j+\frac{N-1}{2}\right)\right]=\frac{1}{2}\left[\log\frac{2\rho}{N}+\phi_{(N-1)/2}\left(\frac{N\beta^{2}}{2\rho}\right)\right]\\ &=\log\beta+\frac{1}{2}\left[\phi_{(N-1)/2}\left(z\right)-\log z\right]\,,\end{split}$ (25)
where $z=N\beta^{2}/(2\rho)$ and
@@ 2.2.3 Singular value decompositions
$\mathrm{Cov}^{X}(t)=\frac{1}{N}\bar{X}(t)^{T}\bar{X}(t)=\frac{1}{N}\bar{X}(0)^{T}V^{T}P(t)^{T}P(t)V\bar{X}(0)=\frac{1}{N}\bar{X}(0)^{T}V^{T}B(t)^{T}\,D(t)^{2}B(t)V\bar{X}(0)\,.$ (26)
@@
The singular value decompositions $P(t)=A(t)D(t)B(t)$ for $t=1,2,3,\dots$ may be chosen so that $B(t)$ converges as $t\to\infty$ to a random orthogonal matrix $B\in O(N-1)$. The limit matrix $B$ depends on the choice only up to multiplication of each row separately by ±$\pm 1$. Up to this row-by-row sign ambiguity, $B$ is distributed according to the Haar measure on $O(N-1)$, and in particular, the first row $B_{1}$ is distributed (up to sign) according to the uniform measure on the sphere $S^{N-2}$. Regardless of the choice, $D(t)^{1/t}$ converges to a deterministic diagonal matrix $D$ with distinct positive diagonal entries satisfying $D_{1,1}>\dots>D_{N-1,N-1}$. Moreover, for $i=1,\dots,N-1$ we have $D_{i,i}=\exp(\lambda_{i})$, where $\lambda_{1}>\lambda_{2}>\dots>\lambda_{N-1}$ are the Lyapunov exponents of the sequence $(P(t))_{t\geq 0}$.
@@
[…]
@@
We apply the above result to our random matrix products $P(t)=M(t)\dots M(1)$. We let $P(t)=A(t)D(t)B(t)$ be a singular value decomposition of $P(t)$, chosen so that in each row of $B(t)$ the leftmost nonzero entry is positive. We must check that the conditions of the theorem are satisfied in this case. We have $k=N-1$ and the support of the distribution of $M(1)$ is the group $\mathrm{GL}_{k}(\mathbb{R})$, so $T_{\mu}=\mathrm{GL}_{k}(\mathbb{R})$. It is straightforward to check that $T_{\mu}$ is $p$-strongly irreducible and $p$-contracting for each $p\in\{1,\dots,k-1\}$. The details are spelled out in [4], Chapter 4, Proposition 2.3, which asserts that if $T_{\mu}$ contains an open subset of $\mathrm{SL}_{k}(\mathbb{R})$ then $T_{\mu}$ is $p$-strongly irreducible and $p$-contracting for each $p\in\{1,\dots,k-1\}$. It remains to check that $\mathbb{E}(\log^{+}\|M(1)\|)<\infty$, which is also straightforward; indeed $\mathbb{E}(\|M(1)\|^{2})<\mathbb{E}(\sum_{i,j}M(1)_{i,j}^{2})<\infty$. To see that the Lyapunov exponents are strictly positive, and hence distinct, we must check that $\mathbb{E}(|\log(|\det M(1)|)|)<\infty$. The quantity $\det M(1)$ is a non-constant polynomial function of the $(N-1)^{2}$ independent non-degenerate Gaussian matrix entries of $M(1)$, and therefore $\det M(1)$ has a distribution with a continuous density on the real line. (In fact, from [15, §2] one may read off an explicit recursive description of the distribution of $\det M(1)$ as a mixture of almost surely non-degenerate Gaussians, with the recursion indexed by the parameter $N$.) In particular the density of $\det M(1)$ is bounded on a neighbourhood of 00, which implies that $\mathbb{E}(|\log^{-}(|\det M(1)|)|)<\infty$, and we also have $\mathbb{E}(\log^{+}(|\det M(1)|))\leq\mathbb{E}((N-1)\log^{+}(\|M(1)\|))<\infty$ because $\|M(1)\|$ is the largest singular value of $M(1)$, while $|\det(M(1))|$ is the product of the singular values.
The entire sequence $\left(P(t)\right)_{t\geq 1}$ is invariant in law under conjugation by any member of the orthogonal group $O(N-1)$. This implies that the limit matrix $B$ must be distributed (up to its row-by-row sign ambiguity) according to Haar measure on $O(N-1)$, since this is the unique right-invariant Borel probability measure on $O(N-1)$. Consequently, up to sign, $B_{1}$ is distributed according to the uniform measure on the unit sphere $S^{N-2}\subset\mathbb{R}^{N-1}$. Finally, the statements about the dependence of $B$ and $D$ upon the choices of polar decompositions $P(t)=A(t)D(t)B(t)$ follow from the fact that almost surely the singular values of $P(t)$ are distinct for each $t\geq 1$. ∎
@@ 2.3 Proof of Theorem 1
We already computed the explicit value of the maximal Lyapunov exponent $\lambda_{1}$ in §2.2.2. By Lemma 5, to prove part (i) of Theorem 1 it suffices to show that $\lambda_{1}<0$ implies that $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ converges almost surely. This follows from Lemma 7 and equation (26). Indeed, if $\lambda_{1}<0$ then $D_{1,1}=\exp(\lambda_{1})<1$ so $D(t)$ converges exponentially fast to 00. Since $B(t)$ also converges, (26) implies that $\mathrm{Cov}^{X}(t)$ almost surely converges exponentially fast to the zero matrix, which implies the convergence of the series $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$.
Part (ii) concerns the critical case where $\lambda_{1}=0$. We saw in Lemma 6 that $\mathrm{Cov}^{X}_{1,1}(t)$ is a random walk with i.i.d. increments distributed as the log of a non-central $\chi^{2}$ random variable, which has positive finite variance and mean 00 in the critical case. This random walk is almost surely unbounded above. This follows, for example, from Theorem 5.1.7 in [16]. In particular $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ almost surely diverges. It then follows from Lemma 5 that $\mu(t)$ almost surely diverges as $t\to\infty$.
Suppose $\lambda_{1}>0$. Let $\mathscr{G}$ be the $\sigma$-algebra generated by $(\overline{X}(t))_{t\geq 0}$. Let $B=\lim_{t\to\infty}B(t)$ be the $\mathscr{G}$-measurable orthogonal matrix described in Lemma 7, where (to make it well-defined) the choice of signs is made so that the first non-zero entry in each row of $B$ is positive. We will study the sequence of dot products $u\cdot X_{i}(t)$. We will use the conditional Lévy-Borel-Cantelli lemma to show that almost surely, conditional on $\mathscr{G}$, we have $|u\cdot X_{i}(t)|\to\infty$ almost surely for each $i=1,\dots,N$. Conditional on $(\mathscr{G},X(t-1))$, the distribution of $\mu(t)$ is Gaussian with mean $\mu(t-1)$ and covariance $\frac{\rho}{N}\mathrm{Cov}^{X}(t-1)$, and the distribution of each $X_{i}(t)$ is Gaussian with mean $\overline{X}_{i}(t)$ and covariance $\frac{\rho}{N}\mathrm{Cov}^{X}(t-1)$. (Indeed, conditional on $\mathscr{G}$ each of $X_{1}(t),\dots,X_{N}(t)$ and $\mu(t)$ determines the others.) It follows that conditional on $(\mathscr{G},X(t-1))$, the distribution of $u\cdot X_{i}(t)$ is Gaussian; we now compute its variance, using equation (26):
@@
By Lemma 7 we see that as $t\to\infty$ the matrix $D(t)/D_{1,1}(t)$ converges almost surely to the $d$ by $d$ matrix whose entries are all 00 except for a $1$ in the top-left corner. Also $B(t)$ converges a.s. to the $\mathscr{G}$-measurable random limit $B$ in $O(N)$. Hence a.s. we have
@@
Let $W$ denote the $\mathscr{G}$-measurable random variable $|u\cdot(B_{1}V\bar{X}(0))|$, which is the absolute value of the sole entry of the $1$ by $1$ matrix $u\bar{X}(0)^{T}V^{T}B_{1}^{T}$. The assumption that $\mathrm{Cov}^{X}(0)$ is a.s. positive definite implies that the column vector $u\bar{X}(0)^{T}V^{T}$ is a.s. nonzero. Recall that the law of $B_{1}$ is absolutely continuous with respect to the uniform measure on the sphere $S^{N-2}$ (up to its sign ambiguity). Since $B$ is independent of $X(0)$, we deduce that $W\neq 0$ almost surely. For any $R>0$ we may bound the Gaussian density by its maximum and integrate over $(-R,R)$ to obtain
@@
The interpretation of Theorem 2 is that there is a random direction in $\mathbb{R}^{d}$ in which the opinions line up in the limit $t\to\infty$. The law of this direction is the unique invariant measure on $P(\mathbb{R}^{d})$ under the action of the group of linear isometries of the quadratic form $\left(\bar{X}(0)^{T}\,\bar{X}(0)\right)^{-1}$. The convergence of the population of opinions to one-dimensionality occurs exponentially with rate $\lambda_{1}-\lambda_{2}$. Later we will show that $\lambda_{1}-\lambda_{2}$ behaves like $1/N$ in the case $\beta=0$, and is at least $c(\alpha,\beta)/N^{2}$ in the general case.
@@ 2.4 Proof of Theorem 2
We will apply a theorem of Guivarc’h and Raugi [10], which is given in [4, Ch. 4, Thm. 1.2]. Let $k\in\{1,\dots,N-1\}$. Recall that in the proof of Lemma 7 we checked that the random matrix product sequence $\left(P(t)\right)_{t\geq 0}$ is $k$-strongly irreducible and $k$-contracting. Guivarc’h and Raugi showed that these conditions imply that there is a unique invariant measure for the action of the random matrices $M(1),M(2),\dots$ on the projective space $\mathbb{P}\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$ We will explain in detail what this means [^5] Although we are simply extending the computation of Newman [20], the details of this argument were not included in Newman’s paper, and we are aware of some papers on Lyapunov exponents in the mathematical physics literature in which similar calculations are performed with incorrect justification..
@@
for each tuple $(i_{1},\dots,i_{k})$ such that $11\leq i_{1}<\dots<i_{k}\leq N-$ 1. To express $\left(Le_{i_{1}}\right)\wedge\dots\wedge\left(Le_{i_{k}}\right)$ in terms of the orthonormal basis (27), one expands the $k$-fold wedge using multilinearity together with the antisymmetry relation, which says that for any permutation $\sigma$ in the symmetric group $S_{k}$, and any $11\leq i_{1}\leq\dots\leq i_{k}\leq N-$ 1 we have
$e_{\sigma(i_{1})}\wedge\dots\wedge e_{\sigma(i_{k})}=\mathrm{sgn}(\sigma)\,e_{i_{1}}\wedge\dots\wedge e_{i_{k}}\,.$
(In particular, any $k$-fold wedge with repetition is 00, which is why such wedges are not included in the basis.) One can check that if $L_{1}$ and $L_{2}$ are two linear maps from $\mathbb{R}^{N-1}$ to $\mathbb{R}^{N-1}$ then $\widehat{(L_{1}L_{2})}=\hat{L}_{1}\hat{L}_{2}$. The induced action of the identity map is the identity, so $\widehat{(L^{-1})}=(\hat{L})^{-1}$. In particular,
@@
Any invertible linear map on $\bigwedge^{k}\mathbb{R}^{N-1}$ descends to a self-map of the projective space $\mathbb{P}\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$. The theorem of Guivarc’h and Raugi mentioned above states that there is a unique invariant probability measure $\pi_{k}$ for this projective action. Let $\widehat{O(N-1)}$ denote the group of transformations of $\mathbb{P}\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$ induced by the action of $O(N-1)$ on $\mathbb{R}^{N-1}$. Since $\pi_{k}$ is unique, it must be invariant with respect to $\widehat{O(N-1)}$, because the matrices $M(t)$ are invariant in law under conjugation by $O(N-1)$ and hence the matrices $\widehat{M(t)}$ are invariant in law under conjugation by the induced orthogonal transformations of $\bigwedge^{p}\mathbb{R}^{N-1}$. Furthermore, the measure $\pi_{k}$ is supported on the closed subvariety $\overline{D}\subseteq\mathbb{P}\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$ which is the image under projectivization of the set $D$ of decomposable vectors in $\bigwedge^{k}\mathbb{R}^{N-1}$. A non-zero vector in $\bigwedge^{k}\mathbb{R}^{N-1}$ is called decomposable if it is of the form $v_{1}\wedge\dots\wedge v_{k}$, where $v_{1},\dots,v_{k}$ are any $k$ linearly independent vectors in $\mathbb{R}^{N-1}$. The action of $\widehat{O(N-1)}$ on $\overline{D}$ is transitive, and this implies that for any two points $x,y\in D$, the distributions of $\frac{\|\hat{M}x\|}{\|x\|}$ and $\frac{\|\hat{M}y\|}{\|y\|}$ coincide. Indeed, there exists some $U\in O(N-1)$ such that $\hat{U}(x)\propto y$, and the law of $M$ is invariant under conjugation by $U$, so the law of $\hat{M}$ is invariant under conjugation by $\hat{U}$. Let $M_{1}^{k}$ denote the $N-1$ by $k$ matrix whose columns are the first $k$ columns of $M$ where $M$ is a matrix distributed like $M(1)$. Applying Furstenberg’s integral formula to compute the maximal Lyapunov exponent of the sequence $\left(\widehat{P(t)}\right)_{t\geq 1}$, we find that
$\lambda_{1}(\alpha,\beta,N)+\dots+\lambda_{k}(\alpha,\beta,N)$ $=$ $\mathbb{E}\left(\log\frac{\|\hat{M}(e_{1}\wedge\dots\wedge e_{k}\|}{\|e_{1}\wedge\dots\wedge e_{k}\|}\right)$
@@
(Note that in the top-dimensional case $k=N-1$ this reduces to the usual determinant formula $\mathbb{E}\log(|\det M|)$ for the sum of all the Lyapunov exponents.) Comparing the above formulae for $k$ and $k-1$ (for any $12\leq k\leq N-$ 1), we obtain

### v2

@@ 2.2.2 The maximal Lyapunov exponent of $P(t)=M(t)\dots M(1)$ for $t\geq 0$
Following Newman [22], note that for our random matrix $M$, for any orthogonal matrix $Q\in O(N-1)$, the law of $Q^{-1}MQ=Q^{T}MQ$ coincides with the law of $M$. (One just has to check that the entries of $Q^{T}MQ$ are pairwise uncorrelated Gaussians with the correct means and variances, a simple calculation.) It follows that the invariant measure $\pi$ is also invariant with respect to the group O(N1)O(N-1). There is a unique such measure on $P\left(\mathbb{R}^{N-1}\right)$. It also follows that the distribution of $\frac{\|Mv\|}{\|v\|}$ does not depend on $v$, so
@@
where $\psi(x)=\frac{\Gamma^{\prime}(x)}{\Gamma(x)}$ is the digamma function. This follows from the statement that each non-central chi-squared random variable is a Poisson-weighted mixture of standard chi-squared random variables. Precisely, if $J\sim\mathrm{Poi}(\kappa/2)$ and, conditional on $J=j$, $Y$ is a $\chi^{2}$ r.v. with $\nu+2j$ degrees of freedom, then $Y\sim\chi^{2}_{\nu}(\kappa)$. See also e.g. the proof of Lemma 10.1 in [20].
@@
[…]
(the last equality holds only when $\beta>0$) where $z=N\beta^{2}/(2\rho)$ and
@@ 2.2.3 Singular value decompositions
$\begin{split}\mathrm{Cov}^{X}(t)&=\frac{1}{N}\bar{X}(t)^{T}\bar{X}(t)=\frac{1}{N}\bar{X}(0)^{T}V^{T}P(t)^{T}P(t)V\bar{X}(0)\\ &=\frac{1}{N}\bar{X}(0)^{T}V^{T}B(t)^{T}\,D(t)^{2}B(t)V\bar{X}(0)\,.\end{split}$ (27)
@@
The singular value decompositions $P(t)=A(t)D(t)B(t)$ for $t=1,2,3,\dots$ may be chosen so that $B(t)$ converges as $t\to\infty$ to a random orthogonal matrix $B\in O(N-1)$. The limit matrix $B$ depends on the choice only up to multiplication of each row separately by $\pm 1$. Up to this row-by-row sign ambiguity, $B$ is distributed according to the Haar measure on O(N1)O(N-1), and in particular, the first row $B_{1}$ is distributed (up to sign) according to the uniform measure on the sphere $\mathbb{S}^{N-2}$. Regardless of the choice, $D(t)^{1/t}$ converges to a deterministic diagonal matrix $D$ with distinct positive diagonal entries satisfying $D_{1,1}>\dots>D_{N-1,N-1}$. Moreover, for $i=1,\dots,N-1$ we have $D_{i,i}=\exp(\lambda_{i})$, where $\lambda_{1}>\lambda_{2}>\dots>\lambda_{N-1}$ are the Lyapunov exponents of the sequence $(P(t))_{t\geq 0}$.
@@
We will use the results of Guivarc’h and Raugi [10] which are also stated in [4], Chapter 4, Theorem 1.2 and Corollary 1.3. The theorem concerns a sequence of random matrix products $S_{n}=Y_{n}\dots Y_{1}$ where $Y_{1},Y_{2},\dots$ are i.i.d. real $k$ by $k$ matrices with common distribution $\mu$, and $Y_{1}$ is almost surely invertible. Denote by $T_{\mu}$ the smallest closed sub-semigroup of $\mathrm{GL}_{k}(\mathbb{R})$ that contains the support of the distribution of $Y_{1}$. For any $p\in\{1,\dots,k-1\}$, $T_{\mu}$ is defined to be $p$-strongly irreducible if there does not exist a finite union $L$ of proper linear subspaces of the exterior power $\bigwedge^{p}\mathbb{R}^{k}$ such that, for all $M\in T_{\mu}$, $\left(\bigwedge^{p}M\right)(L)$ is contained in $L$. Here $\bigwedge^{p}M$ denotes the unique linear endomorphism of $\bigwedge^{p}\mathbb{R}^{k}$ which maps any decomposable vector of the form $v_{1}\wedge\dots\wedge v_{p}$ to $(Mv_{1})\wedge\dots\wedge(Mv_{p})$. Also for any $p\in\{1,\dots,k-1\}$, $T_{\mu}$ is defined to be $p$-contracting if there exists a sequence $\{M_{n}:n\geq 1\}$ in $T_{\mu}$ such that $\|\bigwedge^{p}M_{n}\|^{-1}\bigwedge^{p}M_{n}$ converges to a rank one matrix. Guivarc’h and Raugi proved that if $T_{\mu}$ is $p$-strongly irreducible and $p$-contracting for every $p\in\{1,\dots,k-1\}$ and in addition $\mathbb{E}(\log^{+}\|Y_{1}\|)<\infty$, (where $\|\cdot\|$ denotes the operator norm with respect to the Euclidean norm), then almost surely the following hold:
@@
We apply the above result to our random matrix products $P(t)=M(t)\dots M(1)$. We let $P(t)=A(t)D(t)B(t)$ be a singular value decomposition of $P(t)$, chosen so that in each row of $B(t)$ the leftmost nonzero entry is positive. We must check that the conditions of the theorem are satisfied in this case. We have k=N1k=N-1 and (thinking of $M(1)$ as a $\mathrm{GL}_{k}(\mathbb{R})$-valued random variable) the support of the distribution of $M(1)$ is the whole group $\mathrm{GL}_{k}(\mathbb{R})$, so $T_{\mu}=\mathrm{GL}_{k}(\mathbb{R})$. It is straightforward to check that $T_{\mu}$ is $p$-strongly irreducible and $p$-contracting for each $p\in\{1,\dots,k-1\}$. The details are spelled out in [4], Chapter 4, Proposition 2.3, p. 83, which asserts that if $T_{\mu}$ contains an open subset of $\mathrm{SL}_{k}(\mathbb{R})$ then $T_{\mu}$ is $p$-strongly irreducible and $p$-contracting for each $p\in\{1,\dots,k-1\}$. It remains to check that $\mathbb{E}(\log^{+}\|M(1)\|)<\infty$, which is also straightforward; indeed $\mathbb{E}(\|M(1)\|^{2})<\mathbb{E}(\sum_{i,j}M(1)_{i,j}^{2})<\infty$. To see that the Lyapunov exponents are strictly positive, and hence distinct, we must check that $\mathbb{E}(|\log(|\det M(1)|)|)<\infty$. The quantity $\det M(1)$ is a non-constant polynomial function of the $(N-1)^{2}$ independent non-degenerate Gaussian matrix entries of $M(1)$, and therefore $\det M(1)$ has a distribution with a continuous density on the real line. (In fact, from [15, Section 2] one may read off an explicit recursive description of the distribution of $\det M(1)$ as a mixture of almost surely non-degenerate Gaussians, with the recursion indexed by the parameter $N$.) In particular the density of $\det M(1)$ is bounded on a neighbourhood of 0, which implies that $\mathbb{E}(|\log^{-}(|\det M(1)|)|)<\infty$, and we also have $\mathbb{E}(\log^{+}(|\det M(1)|))\leq\mathbb{E}((N-1)\log^{+}(\|M(1)\|))<\infty$ because $\|M(1)\|$ is the largest singular value of $M(1)$, while $|\det(M(1))|$ is the product of the singular values.
The entire sequence $\left(P(t)\right)_{t\geq 1}$ is invariant in law under conjugation by any member of the orthogonal group O(N1)O(N-1). This implies that the limit matrix $B$ must be distributed (up to its row-by-row sign ambiguity) according to Haar measure on O(N1)O(N-1), since this is the unique right-invariant Borel probability measure on O(N1)O(N-1). Consequently, up to sign, $B_{1}$ is distributed according to the uniform measure on the unit sphere $\mathbb{S}^{N-2}\subset\mathbb{R}^{N-1}$. Finally, the statements about the dependence of $B$ and $D$ upon the choices of polar decompositions $P(t)=A(t)D(t)B(t)$ follow from the fact that almost surely the singular values of $P(t)$ are distinct for each $t\geq 1$. ∎
@@ 2.3 Proof of Theorem 1
We already computed the explicit value of the maximal Lyapunov exponent $\lambda_{1}$ in Section 2.2.2. By Lemma 5, to prove part (i) of Theorem 1 it suffices to show that $\lambda_{1}<0$ implies that $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ converges almost surely. This follows from Lemma 7 and equation (27). Indeed, if $\lambda_{1}<0$ then $D_{1,1}=\exp(\lambda_{1})<1$ so $D(t)$ converges exponentially fast to 0. Since $B(t)$ also converges, (27) implies that $\mathrm{Cov}^{X}(t)$ almost surely converges exponentially fast to the zero matrix, which implies the convergence of the series $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$.
Part (ii) concerns the critical case where $\lambda_{1}=0$. We saw in Lemma 6 that $\mathrm{Cov}^{X}_{1,1}(t)$ is a random walk with i.i.d. increments distributed as the log of a non-central $\chi^{2}$ random variable, which has positive finite variance and mean 0 in the critical case. This random walk is almost surely unbounded above. This follows, for example, from Theorem 5.1.7 in [18]. In particular $\sum_{t=0}^{\infty}\mathrm{Cov}^{X}(t)$ almost surely diverges. It then follows from Lemma 5 that $\mu(t)$ almost surely diverges as $t\to\infty$.
Suppose $\lambda_{1}>0$. Let $\mathscr{G}$ be the $\sigma$-algebra generated by $(\overline{X}(t))_{t\geq 0}$. Let $B=\lim_{t\to\infty}B(t)$ be the $\mathscr{G}$-measurable orthogonal matrix described in Lemma 7, where (to make it well-defined) the choice of signs is made so that the first non-zero entry in each row of $B$ is positive. We will study the sequence of dot products $u\cdot X_{i}(t)$. We will use the conditional Lévy-Borel-Cantelli lemma to show that almost surely, conditional on $\mathscr{G}$, we have $|u\cdot X_{i}(t)|\to\infty$ almost surely for each $i=1,\dots,N$. Conditional on $(\mathscr{G},X(t-1))$, the distribution of $\mu(t)$ is Gaussian with mean $\mu(t-1)$ and covariance $\frac{\rho}{N}\mathrm{Cov}^{X}(t-1)$, and the distribution of each $X_{i}(t)$ is Gaussian with mean $\overline{X}_{i}(t)$ and covariance $\frac{\rho}{N}\mathrm{Cov}^{X}(t-1)$. (Indeed, conditional on $\mathscr{G}$ each of $X_{1}(t),\dots,X_{N}(t)$ and $\mu(t)$ determines the others.) It follows that conditional on $(\mathscr{G},X(t-1))$, the distribution of $u\cdot X_{i}(t)$ is Gaussian. We now compute its variance using equation (27):
@@
By Lemma 7 we have $D_{i,i}(t)^{1/t}\to\exp(\lambda_{i})$ as $t\to\infty$, so the matrix $D(t)/D_{1,1}(t)$ converges almost surely to the N1N-1 by N1N-1 matrix whose entries are all 0 except for a 11 in the top-left corner. Also $B(t)$ converges a.s. to the $\mathscr{G}$-measurable random limit $B$ in $O(N)$; recall that $B_{1}$ is the first row of $B$. Hence a.s. we have
@@
Let $W$ denote the $\mathscr{G}$-measurable random variable $|u\cdot(B_{1}V\bar{X}(0))|$, which is the absolute value of the sole entry of the $1$ by $1$ matrix $u\bar{X}(0)^{T}V^{T}B_{1}^{T}$. The assumption that $\mathrm{Cov}^{X}(0)$ is a.s. positive definite implies that the column vector $u\bar{X}(0)^{T}V^{T}$ is a.s. nonzero. Recall that the law of $B_{1}$ is absolutely continuous with respect to the uniform measure on the sphere $\mathbb{S}^{N-2}$ (up to its sign ambiguity). Since $B$ is independent of $X(0)$, we deduce that $W\neq 0$ almost surely. For any $R>0$ we may bound the Gaussian density by its maximum and integrate over (R,R)(-R,R) to obtain
@@ 2.4 Proof of Theorem 2
We will apply a theorem of Guivarc’h and Raugi [10], which is given in [4, Ch. 4, Thm. 1.2]. Let $k\in\{1,\dots,N-1\}$. Recall that in the proof of Lemma 7 we checked that the random matrix product sequence $\left(P(t)\right)_{t\geq 0}$ is $k$-strongly irreducible and $k$-contracting. Guivarc’h and Raugi showed that these conditions imply that there is a unique invariant measure for the action of the random matrices $M(1),M(2),\dots$ on the projective space $P\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$ . We will explain in detail what this means [^4] Although we are simply extending the computation of Newman [22], the details of this argument were not included in Newman’s paper, and we are aware of some papers on Lyapunov exponents in the mathematical physics literature in which similar calculations are performed with incorrect justification..
@@
for each tuple $(i_{1},\dots,i_{k})$ such that $1\leq i_{1}<\dots<i_{k}\leq N-1$.
To express $\left(Le_{i_{1}}\right)\wedge\dots\wedge\left(Le_{i_{k}}\right)$ in terms of the orthonormal basis (28), one expands the $k$-fold wedge using multilinearity together with the antisymmetry relation, which says that for any permutation $\sigma$ in the symmetric group $S_{k}$, and any $1\leq i_{1}\leq\dots\leq i_{k}\leq N-1$ we have
$e_{i_{\sigma(1)}}\wedge\dots\wedge e_{i_{\sigma(k)}}=\mathrm{sgn}(\sigma)\,e_{i_{1}}\wedge\dots\wedge e_{i_{k}}\,.$
(In particular, any $k$-fold wedge with repetition is 0, which is why such wedges are not included in the basis.) One can check that if $L_{1}$ and $L_{2}$ are two linear maps from $\mathbb{R}^{N-1}$ to $\mathbb{R}^{N-1}$ then $\widehat{(L_{1}L_{2})}=\hat{L}_{1}\hat{L}_{2}$. The induced action of the identity map is the identity, so $\widehat{(L^{-1})}=(\hat{L})^{-1}$. In particular,
@@
Any invertible linear map on $\bigwedge^{k}\mathbb{R}^{N-1}$ descends to a self-map of the projective space $P\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$. The theorem of Guivarc’h and Raugi mentioned above states that there is a unique invariant probability measure $\pi_{k}$ for this projective action. Let $\widehat{O(N-1)}$ denote the group of transformations of $P\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$ induced by the action of O(N1)O(N-1) on $\mathbb{R}^{N-1}$. Since $\pi_{k}$ is unique, it must be invariant with respect to $\widehat{O(N-1)}$, because the matrices $M(t)$ are invariant in law under conjugation by O(N1)O(N-1) and hence the matrices $\widehat{M(t)}$ are invariant in law under conjugation by the induced orthogonal transformations of $\bigwedge^{p}\mathbb{R}^{N-1}$. Furthermore, the measure $\pi_{k}$ is supported on the closed subvariety $\overline{D}\subseteq P\left(\bigwedge^{k}\mathbb{R}^{N-1}\right)$ which is the image under projectivization of the set $D$ of decomposable vectors in $\bigwedge^{k}\mathbb{R}^{N-1}$. A non-zero vector in $\bigwedge^{k}\mathbb{R}^{N-1}$ is called decomposable if it is of the form $v_{1}\wedge\dots\wedge v_{k}$, where $v_{1},\dots,v_{k}$ are any $k$ linearly independent vectors in $\mathbb{R}^{N-1}$.
Let $M$ be a random matrix distributed like $M(1)$, and let $\hat{M}$ be the map that $M$ induces on $\bigwedge^{k}\mathbb{R}^{N-1}$. The action of $\widehat{O(N-1)}$ on $\overline{D}$ is transitive, and this implies that for any two points $x,y\in D$, the distributions of $\frac{\|\hat{M}x\|}{\|x\|}$ and $\frac{\|\hat{M}y\|}{\|y\|}$ coincide. Indeed, there exists some $U\in O(N-1)$ such that $\hat{U}(x)\propto y$, and the law of $M$ is invariant under conjugation by $U$, so the law of $\hat{M}$ is invariant under conjugation by $\hat{U}$.
Let $M_{1}^{k}$ denote the N1N-1 by $k$ matrix whose columns are the first $k$ columns of $M$. Applying Furstenberg’s integral formula to compute the maximal Lyapunov exponent of the sequence $\left(\widehat{P(t)}\right)_{t\geq 1}$, we find that
$\lambda_{1}(\alpha,\beta,N)+\dots+\lambda_{k}(\alpha,\beta,N)$ $=$ $\mathbb{E}\left(\log\frac{\|\hat{M}(e_{1}\wedge\dots\wedge e_{k})\|}{\|e_{1}\wedge\dots\wedge e_{k}\|}\right)$
@@
(Note that in the top-dimensional case k=N1k=N-1 this reduces to the usual determinant formula $\mathbb{E}\log(|\det M|)$ for the sum of all the Lyapunov exponents.) Comparing the above formulae for $k$ and k1k-1 (for any $2\leq k\leq N-1$), we obtain